0.1m3 of gas is compressed from pressure of 120KN/m2 and temperature 250C to a pressure of 1.2MN/maccording to the law PV1.2 = constant.

Determine,

(a) The work transferred during the compression,

(b) The change in internal energy,

(c) The heat transferred during the compression

Assume , cv=0.72kJ/kgK         and           R=0.285KJ/kgK


1
Expert's answer
2022-03-10T10:49:17-0500

cv=0.72kJ/kgK


R=0.285KJ/kgK


a.) "work = \u2212 P \u0394 V"


For PV"^{\gamma}" = constant


work done by the gas = "\frac{(P_2V_2-P_1V_1)}{(1-\gamma)}" where "\gamma = 1.2"

for calculation of "P_2 = P_1(\frac{V_1}{V_2})^{1.2}"


"= 1.2(\frac{1.2}{120})^{120}"


P2= 3.175


hence work done="\frac{3.175\times1\times0.05}{1.2-1}"


work done by gas = 0.794 m3


b.) "\u0394U = Q \u2212 W"


"\u0394U =" 18J - 7.125J


= 10.875 J


c.) "\u0394H=qp"


= "\frac{25\times1.2}{120}"


= "\frac{30}{120}"


= 0.25